OpenStudy (anonymous):

3x is a factor of 3x^2-27x, what is the other factor

OpenStudy (skullpatrol):

$$\huge 3x^2-27x$$

OpenStudy (kc_kennylau):

\[3x^2-27x\]\(=x(3x-27)\)\[=3x(x-9)\]

OpenStudy (kc_kennylau):

So we know that \(3x\) is one of the factors of \(3x^2-27x\). \(3x^2\) is \(x\) times \(3x\) and \(27x\) is \(9\) times \(3x\). So we can express \(3x^2-27x\) as \(x\times3x-9\times3x\). Take out the common factor and we have \(3x(x-9)\). Therefore its other factor is \(x-9\). :)

OpenStudy (anonymous):

yaaas thankyou

OpenStudy (kc_kennylau):

no problem :)